# Robust Offline Imitation Learning Through State-level Trajectory Stitching

### v1

@@ Abstract
Imitation learning (IL) has proven effective for enabling robots to acquire visuomotor skills through expert demonstrations. However, traditional IL methods are limited by their reliance on high-quality, often scarce, expert data, and suffer from covariate shift. To address these challenges, recent advances in offline IL have incorporated suboptimal, unlabeled datasets into the training. In this paper, we propose a novel approach to enhance policy learning from mixed-quality offline datasets by leveraging task-relevant trajectory fragments and rich environmental dynamics. Specifically, we introduce a state-based search framework that stitches state-action pairs from imperfect demonstrations, generating more diverse and informative training trajectories. Experimental results on standard IL benchmarks and real-world robotic tasks showcase that our proposed method significantly improves both generalization and performance. Videos are available on https://youtube.com/sbr.
@@ I Introduction
Manually collected data often suffer from behavioral inconsistencies due to varying operator skills, making it difficult to obtain reliable expert demonstrations. Moreover, these demonstrations may include suboptimal actions, such as retries of failed grasps [11]. Directly applying behavioral cloning [12] to such data can lead to policies that replicate undesirable failure behaviors [13]. To overcome these limitations, algorithms must be capable of identifying and utilizing beneficial segments from imperfect demonstrations while filtering out irrelevant or noisy actions. This process, often referred to as trajectory stitching in reinforcement learning [13], is crucial for improving the robustness of the policy learned from imperfect data.
@@
Recent research has addressed the problem of learning from mixed-quality datasets, where offline data include both expert data and suboptimal, unlabeled demonstrations [14, 15, 16, 17, 18, 19, 20, 11, 21]. These imperfect demonstrations often originate from failed attempts or random policies or even from different tasks executed within the same environment [22]. Despite not perfectly matching expert behavior, such data can provide valuable insights—especially in their early stages. For example, in the task of pouring water into the purple cup (see Figure 1), a failed demonstration involving pouring water into the blue cup still shares crucial behavioral patterns with the expert’s actions in the early stages. Furthermore, imperfect trajectories may include states outside the expert distribution, which can offer rich transitional information from non-expert to expert states. These transitions teach the agent how to handle deviations from the expert’s state distribution. However, such valuable information is often overlooked in many existing IL frameworks.
@@ III Policy Learning via State-based Search and Trajectory Stitching
Figure 2: Overview of the proposed framework: (a) Representation Learning utilizes a pre-trained representation model to encode states, establishing a similarity metric based on latent space distances. (b) State-based Search and Extraction identifies trajectory segments reflecting beneficial behavior through state-based search, augmenting the training set. (c) Policy Training fine-tunes an IL policy on the union of expert and retrieved data.
In this section, we present our methodology for learning policies from datasets of varying quality in offline IL. Our framework consists of several key stages, outlined as follows: First, we formally define the problem in Section III-A. Next, we introduce a state similarity metric based on a well-trained encoder in Section III-B. This is followed by the state-based search process, described in Section III-C, which retrieves and stitches together useful trajectory segments from both expert and imperfect data. Finally, in Section III-D, we detail the policy training phase, where we fine-tune the IL policy on the union of expert and retrieved data. A schematic representation of the framework is shown in Figure 2.
@@ III-C State-based Search and Extraction
$\mathcal{D}_{u}^{0}\leftarrow\mathcal{D}_{u}^{0}\cup\left\{s_{t-1}^{o}\big{|}\mathcal{F}_{\mathcal{D}_{e}}(s_{t}^{o})>\delta,\,t-1\geq 0,\,s_{t}^{o}\in\mathcal{D}_{o}\right\}.$ (5)
@@
$\mathcal{D}_{u}^{k}$ $\leftarrow\mathcal{D}_{u}^{k}\cup\left\{s_{t-1}^{o}\big{|}\mathcal{F}_{D^{k}_{e}}(s_{t}^{o})\!>\!\delta,\,t-\!1\geq\!0,\,s_{t}^{o}\in\!D^{k}_{o}\right\},$ (7)
@@ III-D Behavior Cloning
5:     Sample $s\sim\mathcal{D}_{e}\cup\mathcal{D}_{o}$
@@ IV Experiments
TABLE I: Performance comparison on the D4RL benchmark. All results are normalized between 00 and $100$, where 00 indicates random policy performance, and $100$ corresponds to expert-level performance [22]. The best performance in each category is highlighted in bold, while second- and third-place results are underlined.
@@
door-cloned-v1 $\underline{5.25\pm 7.90}$ $0.02\pm 0.04$ $0.07\pm 0.09$ $0.36\pm 0.33$$-.33\pm 0.01$ $\underline{8.28\pm 7.82}$ $\mathbf{25.6\pm 15.9}$
@@
relocate-cloned-v1 $\underline{59.8\pm 32.9}$ $10.5\pm 5.23$ $-0.1\pm 0.09$ $56.2\pm 23.7$$-.32\pm 0.03$ $\underline{61.6\pm 3.57}$ $\mathbf{69.9\pm 6.38}$
@@ IV-A2 Real-world Environments
For real-world data collection, we deploy a teleoperation system where the operator controls the robot via a 3Dconnexion SpaceMouse at $10$Hz during demonstrations. For each task, we collect only $20$ demonstrations as expert data, which is significantly fewer than in previous works [2, 3, 4, 39]. The imperfect demonstrations come from approximately $100$ failed trajectories during the collection process, as well as trajectories with different subtask goals or random arm movements. The UR5 robot station is equipped with two RealSense D435 depth cameras, which are downsampled to a 256x256 resolution and used for policy input. Additionally, the policy incorporates the robot arm’s proprioception as input. For all tasks, we employ position control, with the policy directly outputting the desired end-effector pose.
@@ IV-B Baselines
For simulation tasks, we use a three-layer perceptron as the policy network. In real-world robotic tasks, we employ BAKU, a transformer-based policy network. Additionally, to address the challenge of data scarcity, we incorporate R3M [40], a pre-trained feature representation for robotic manipulation tasks.
@@ IV-D Ablation Study
Figure 6: Results of real-world experiments. The policy success rates averaged over $20$ trials for each setup are reported.

### v2

@@ Abstract
Imitation learning (IL) has proven effective for enabling robots to acquire visuomotor skills through expert demonstrations. However, traditional IL methods are limited by their reliance on high-quality, often scarce, expert data, and suffer from covariate shift. To address these challenges, recent advances in offline IL have incorporated suboptimal, unlabeled datasets into the training. In this paper, we propose a novel approach to enhance policy learning from mixed-quality offline datasets by leveraging task-relevant trajectory fragments and rich environmental dynamics. Specifically, we introduce a state-based search framework that stitches state-action pairs from imperfect demonstrations, generating more diverse and informative training trajectories. Experimental results on standard IL benchmarks and real-world robotic tasks showcase that our proposed method significantly improves both generalization and performance. Videos are available on https://www.youtube.com/playlist?list=PLUzaZQK4aUT3SER8YY-T_dBnE_ZS9_9_w.
@@ I Introduction
Manually collected data often suffer from behavioral inconsistencies due to varying operator skills, making it difficult to obtain reliable expert demonstrations. Moreover, these demonstrations may include suboptimal actions, such as retries of failed grasps [11]. Directly applying behavioral cloning [12] to such data can lead to policies that replicate undesirable failure behaviors [13]. To overcome these limitations, algorithms must be capable of identifying and utilizing beneficial segments from imperfect demonstrations while filtering out certain irrelevant or noisy actions. This process, often referred to as trajectory stitching in reinforcement learning [13], is crucial for improving the robustness of the policy learned from imperfect data.
@@
Recent research has addressed the problem of learning from mixed-quality datasets, where offline data include both expert data and suboptimal, unlabeled demonstrations [14, 15, 16, 17, 18, 19, 20, 11, 21]. These imperfect demonstrations often originate from failed attempts or random policies or even from different tasks executed within the same environment [22]. Despite not perfectly matching expert behavior, such data can provide valuable insights. For example, in the task of pouring water into the purple cup (see Figure 1), a failed demonstration involving pouring water into the blue cup still shares crucial behavioral patterns with the expert’s actions in the early stages. Furthermore, imperfect trajectories may include states outside the expert distribution, which can offer rich transitional information from non-expert to expert states. These transitions teach the agent how to handle deviations from the expert’s state distribution. However, such valuable information is often overlooked in many existing IL frameworks.
@@ III Policy Learning via State-based Search and Trajectory Stitching
Figure 2: Overview of the proposed framework: (a) Representation Learning utilizes a pre-trained representation model to encode states, establishing a similarity metric based on latent space distances. (b) State-based Search and Extraction identifies trajectory segments reflecting beneficial behavior through state-based search, augmenting the training set. (c) Policy Training trains an IL policy on the union of expert and retrieved data.
In this section, we present our methodology for learning policies from datasets of varying quality in offline IL. Our framework consists of several key stages, outlined as follows: First, we formally define the problem in Section III-A. Next, we introduce a state similarity metric based on a well-trained encoder in Section III-B. This is followed by the state-based search process, described in Section III-C, which retrieves and stitches together useful trajectory segments from imperfect data. Finally, in Section III-D, we detail the policy training phase, where we train the IL policy on the union of expert and retrieved data. A schematic representation of the framework is shown in Figure 2.
@@ III-C State-based Search and Extraction
$\mathcal{D}_{u}^{0}\leftarrow\left\{s_{t-1}^{o}\big{|}\mathcal{F}_{\mathcal{D}_{e}}(s_{t}^{o})>\delta,\,t-1\geq 0,\,s_{t}^{o}\in\mathcal{D}_{o}\right\}.$ (5)
@@
$\mathcal{D}_{u}^{k}$ $\leftarrow\left\{s_{t-1}^{o}\big{|}\mathcal{F}_{D^{k}_{e}}(s_{t}^{o})\!>\!\delta,\,t-\!1\geq\!0,\,s_{t}^{o}\in\!D^{k}_{o}\right\},$ (7)
@@ III-D Behavior Cloning
5:  Sample $(s_{t},a_{t},s_{t+1})\sim\mathcal{D}_{e}\cup\mathcal{D}_{o}$
@@ IV Experiments
TABLE I: Performance comparison on the D4RL benchmark. All results are normalized between 0 and $100$, where 0 indicates random policy performance, and $100$ corresponds to expert-level performance [22]. The best performance in each category is highlighted in bold, while second- and third-place results are underlined.
@@
door-cloned-v1 $\underline{5.25\pm 7.90}$ $02\pm 0.04 0.0$$07\pm 0.09$ $6\pm 0.33 .33±0.0$1-$3\pm$ 0.01 $\underline{8.28\pm 7.82}$ $\mathbf{25.6\pm 15.9}$
@@
relocate-cloned-v1 $\underline{59.8\pm 32.9}$ $10.5\pm 5.23$ $0.1\pm 0.09$ $56.2\pm 23.7$ $32\pm 0.03$ $\underline{61.6\pm 3.57}$ $\mathbf{69.9\pm 6.38}$
@@ IV-A2 Real-world Environments
For real-world data collection, we deploy a teleoperation system where the operator controls the robot via a 3Dconnexion SpaceMouse at 1010Hz during demonstrations. For each task, we collect only $20$ demonstrations as expert data, which is significantly fewer than in previous works [2, 3, 4, 39]. The imperfect demonstrations including image observations come from approximately $100$ failed trajectories during the collection process, as well as trajectories with different subtask goals or random arm movements. The UR5 robot station is equipped with two RealSense D435 depth cameras, which are downsampled to a 256x256 resolution and used for policy input. Additionally, the policy incorporates the robot arm’s proprioception as input. For all tasks, we employ position control, with the policy directly outputting the desired end-effector pose.
@@ IV-B Baselines
For simulation tasks, we use a three-layer perceptron as the policy network. In real-world robotic tasks, we employ BAKU [1], a transformer-based policy network. Additionally, to calculate the similarity between pixel states, we incorporate R3M [40], a pre-trained feature representation for robotic manipulation tasks.
@@ IV-D Ablation Study
Figure 6: Results of real-world experiments. The policy success rates averaged over 2020 trials for each setup are reported.